ZL = 200 Ω and it is desired that Zi = 50 Ω The quarter wave transformer should have a characteristic impedance of


A. 100 Ω

B. 40 Ω

C. 10000 Ω

D. 4 Ω


1
Expert's answer
2021-01-07T11:00:14-0500

Solution:

"Z_O= \sqrt{\smash[b]{Z_I\cdot Z_L}}= \sqrt{\smash[b]{200\cdot50}}=100"


"Z_L=Load \;impidence"

"Z_I=Input\;impidence"

"Z_O=Characteristic \;impidece"


Answer:

"Characteristic \;impidece\;is\;100\Omega"